The Right Way to Tax DAT

Richard M. Stallman

Record company magnates don’t like the digital audio tape recorder (DAT), which can make perfect copies of musical recordings. They fear that customers will copy music themselves, and stop buying prerecorded music.

Threatening lawsuits, they have obtained from the manufacturers of DATs an agreement to pay a fee for each DAT unit and each DAT tape sold to consumers. This fee is to be divided among various participants in the music business: musicians, composers, music publishers and record companies. In addition, DAT manufacturers have agreed to cripple DAT units so that they cannot make a copy of a copy of a prerecorded piece.

Now the record companies have asked Congress to enact a law turning this fee into a tax and prohibiting manufacture of DAT tapedecks that function without imposed limitations.

The stated purpose of the tax is to “compensate” musicians for copying done by individuals using DATs. However, 57 percent of the funds collected would go to record companies and music publishers—leaving less than half to the people who participate in the creative process. Most of these remaining funds would go to musical superstars, and thus would do little to encourage musical creativity. Meanwhile, DAT users would be unable to make full use of the power of DAT technology.

Here is a proposal for a different system for taxing DATs and DAT tape—one designed to support music rather than cater to vested interests.

• Collect funds with a tax on DAT machines and DAT tapes, as the current proposal provides.

• Use a survey system to measure the extent of copying of each musical piece.

• Distribute these funds entirely to the people who create music.

• Adjust each contributor’s share so that it increases more slowly per copy as it gets larger. This spreads the funds more widely to support a larger number of musicians adequately.

• Make no restrictions on the functioning of DATs.

What is the purpose of copyright?

The record industry presents its proposal as a way to “compensate” musicians, assuming that they are entitled to be paid for any copy made. Many Americans believe that copyright law reflects a natural right of authors or musicians—that these are entitled to special consideration from public policy. However, any lawyer specializing in the field knows this is a misunderstanding, a view rejected by the American legal system.

The stated purpose of copyright, given in the U.S. Constitution, is to “promote the progress of science and the useful arts”. Progress in music means new and varied music for the public to enjoy: copyright is supposed to promote a public good, not a private one.

Yet copyright is often thought of as a natural right by laymen and politicians, which often leads to wrong decisions about copyright policy. Even courts, defining the details of the copyright system, often let this thought creep back implicitly even though it is supposed to be excluded. This is a conceptual error because it mistakes a means (copyright) to a larger end (progress) for an end in itself.

Promoting progress in the arts does not inherently justify the idea that authors are entitled to any particular sort of copyright, or even that copyright should exist at all. Copyright is justified if the benefits of progress exceeds the burden that copyright imposes on everyone except the copyright holder.

How do we make this cost/benefit comparison? It depends partly on facts (how does a particular law affect musical activity and music users) and partly on our value judgements about those results.

Let’s assume that it is worth paying a DAT tax if the result is a significant increase in musical activity, and investigate how we should arrange the details of this tax in order to maximize the benefit. But first, let’s review basic principles and facts which have a bearing on the inquiry.

Diminishing returns

The law of diminishing returns is a general principle of economics. It states that each additional increment of efforts or funds spent on a given goal typically produces a smaller and smaller increment in the results. There are exceptions to this law, but they are local; if you keep on increasing the inputs, you eventually leave the exceptions behind.

For example, you can make traffic flow more smoothly by improving roads. Adding one lane to 20 miles of congested roads in a city might increase the average traffic speed by 15 miles an hour. Adding a second lane to those roads will not give the same improvement; this might increase the average speed by only 5 more miles an hour. The next additional lane might make no noticeable difference if the traffic jams are already gone. Yet each successive lane will cause greater dislocation as more and more buildings must be torn down to make room.

When applied to the activities of musicians, diminishing returns tells us that each successive increase in the income of musicians will have a smaller effect on the amount of creativity in music.

Diminishing returns is the first reason to reject the idea that any use of music “should” be covered by copyright. There is nothing to gain by trying to guarantee owners control of every possible aspect of the use of music or to give them a financial stake in every possible aftermarket. Extending copyright can only “promote progress” up to a certain point. Further extensions merely increase what the public pays to the owners for what they will do anyway. Extending copyright beyond that point is certainly undesirable.

Trade-offs

Those with a vested interest in extending copyright start the discussion by claiming that copyright “should” be extended as far as it can go. But the principle of diminishing returns renders this claim implausible. So they fall back on the position that copyright should be extended to maximize the rate of progress. But this too is wrong, because it ignores the existence of other trade-offs. Copyright imposes costs and burdens on the public, like any other government project. The benefit may not be worth the price.

Government fills many important functions, but few would say that any one of these functions should be expanded to maximize output. For example, governments build roads, and this is very useful. But few leaders would advocate building every road that could be built. Road construction is expensive, and citizens have other uses for their money. Too much concentration on building roads means that other social and individual needs will be unmet.

The same considerations apply to individual decisions. By spending more money, you can buy a bigger and fancier house. Most people would prefer the more expensive house, all else being equal. But given finite resources, at some point spending more on a house becomes a poor allocation of them.

Copyright does not directly spend public funds, but it does impose a cost—a loss of freedom— on every citizen. The wider the scope of copyright, the more freedom we pay. We might prefer to exercise some of our freedoms rather than trade them away. We must judge any decision in copyright policy by comparing the benefits with the costs.

Incentive is the wrong concept

The idea of providing a monetary incentive for making music is based on a misunderstanding. Musicians hope primarily for other kinds of reward; they must. Very few musicians get rich from their music; a talented person whose primary goal is wealth would seek it in other ways.

In fact, psychological studies show that the desire for an extrinsic reward (such as profit) generally hampers creative activities such as writing music. The people who can do them well are usually those who do them mostly for their own sake.

This is not to say that musicians don’t care about being paid. Most hope to make a living from music so they will be free to devote their time to it. As long as they earn enough to live, they will make music as best they can. We might wish them to earn somewhat more than just enough, so they can live as well as most Americans. But to offer them wealth beyond this gains the public little—it is a matter of diminishing returns.

With this understanding, let’s consider how a tax on DAT tape could be designed to serve the intended purpose of copyright.

Who should get the funds

If the purpose of the DAT tax is to better reward musicians and composers, then all the money collected should go to them—not just 43 percent. The musicians and composers are the ones who truly create the music. In principle, we could do without record companies entirely.

Record companies do provide a useful service: they distribute prerecorded copies of music, usually of high quality. This service is widely used, and will probably remain so. And it is right that the purchasers of prerecorded copies should pay for this service. But listeners making copies for themselves or their friends do not consume this service; they use only the work of the musicians and composers. The record companies contribute only incidentally and their role is not essential.

Dividing the funds

What share of the tax revenues should each musician or composer get? The record company proposal would divide the money in proportion to record sales.

It makes sense to distribute the funds based on how much that musician’s work is copied, more or less. But strict proportionality is not the best apportionment. If each musician gets a share in strict proportion to the amount of copying of his or her music, then a large share will go to make a few superstars even richer than they are now. This won’t do much to promote musical culture or diversity.

We can promote music more effectively by making any one musician’s share of the tax revenues taper off as copies increase. For example, we could calculate an “adjusted number of copies” which, beyond a certain point, increases more slowly than the actual number.

The effect of tapering off will be to spread the money more widely, supporting more musicians at an adequate standard of living. This encourages diversity, which is what copyright is supposed to do.

The US government has already established a program to fund diversity in the arts: the NEA. However, NEA grants involve discretionary power, which makes them a center for controversy, sometimes because a few members of the public strongly dislike the work, and sometimes because hardly anyone particularly likes it. Spreading out DAT tax revenues will also have the effect of supporting less popular musicians. However, it will not support musicians whose work nobody likes. In addition, since it involves no discretion, no arbitrary decisions, there is little room for objection on account of any particular case.

Encouraging home copying

The record company proposal includes a requirement to make it difficult for home listeners to make copies. Specifically, it requires that consumer DAT machines refuse to copy a copy that was made on a consumer DAT machine. The argument for this requirement is based on the assumption that home copying is somehow unfair.

In the past, many people have considered it unfair, because it reduced the income of musicians. The DAT tax makes this reason obsolete. Once home copying does contribute to the income of musicians, through the DAT tax, the reason to discourage home copying disappears.

Therefore, if a DAT tax is adopted, the ability to copy DAT tapes should not be restricted. Home copying is more efficient than record companies and record stores; music lovers should be encouraged to use home copying as much as possible.

Measuring the use of each piece of music

Today, nearly all the recorded music in the United States is purchased in record stores; home copying is but a small fraction. This will probably remain true for a long time, because record stores offer a place where a person can go to find a particular piece or to browse a wide selection. While this remains true, we can usually estimate the audience of a given piece fairly well by counting record sales.

Eventually, home copying may become so widespread that estimating its extent from sales figures may be unsatisfactory. This is already unsatisfactory for musicians who distribute independently without the help of record companies; and if any musicians need additional support, these are the ones. We need another way to estimate usage of any given piece, in order to distribute the tax funds.

We can make these estimates by survey. From time to time, survey staff would ask randomly chosen members of the public to show what copies they have made of copyrighted music. The citizens asked would not be required to answer. But no penalty and no guilt would attach to having made copies, so most people will be glad to participate. Fans will hope to be chosen so that they can contribute to the count for their favorite musical groups.

To make the survey more efficient and broader-based (and thus more accurate), it could be automated. The survey bureau could mail read-write memory cards to the chosen participants, who would connect them momentarily to their DAT units and then mail them back. With proper design, the survey bureau would have no way of knowing who had sent in any particular card, and thus no information about who had copied what, but they would still have an accurate total.

Conclusion

The record companies have proposed an excellent scheme for taxing the public to increase their own income, but this isn’t a legitimate purpose of copyright. Through due attention to the ends of copyright rather than past means, we can design a system which supports musicians while giving citizens full freedom to copy music as they wish.

What You Can Do

Record company lobbyists are working hard to pass their form of DAT tax. There is little organized opposition, and little public debate. Their bill has already been sent out of committee in the Senate.

This article proposes an alternative to the record company plan. In order for this alternative, or any alternative, to have a chance, we must first prevent the hasty adoption of the record company plan. To help accomplish this, please write letters to:

Congressman Barney Frank 437 Cherry St West Newton, MA 02165

Senator Metzenbaum United States Senate Washington, DC 20510

House Subcommittee on Intellectual Property House of Representatives Washington, DC 20515

Urge Congress to reject the record company bill so that this and other alternatives can be properly considered. It takes just a few minutes to write a short letter, but in combination with other people’s letters it can do a great deal of good.

If you know any musicians, composers, or songwriters, give them copies of this article. Many musicians prefer this alternative to the record company tax plan, and they are strongly motivated to act on their concern.